OpenStudy (anonymous):

Jeremy ran 440 meters each day for 5 days during physical education class. How many total kilometers did he run during these 5 days? F. 2.2 G. 22 H. 220 J. 2200

OpenStudy (saifoo.khan):

440*5 2200

OpenStudy (anonymous):

440*5=2200

OpenStudy (anonymous):

You have to convert 2200 into kilmeters

OpenStudy (anonymous):

\[\frac{440meter}{day} \cdot 5day = (440\cdot 5) meter\]

OpenStudy (anonymous):

\[(440\cdot 5) meter \cdot \frac{1kilometer}{1000meter} = \frac{440\cdot5}{1000}kilometer\]

OpenStudy (saifoo.khan):

OIC, sorry my bad!

OpenStudy (saifoo.khan):

0.44 x 5 equals?

OpenStudy (anonymous):

2.2

OpenStudy (saifoo.khan):

\[\Huge O-yea!.. 2.2\]